Prove that the plane 2x-3y+6z=6 touches the conicoid 4x²-9y²+36z²=36
Find the point of contact.
1
Expert's answer
2019-07-12T10:03:24-0400

1. Represent the plain and the conicoid in an implicit form:

"F(x,y,z)=2x-3y+6z-6=0;"

"G(x,y,z)=4x\u00b2-9y\u00b2+36z\u00b2-36=0."


2. The normal vectors to the surfaces are determined by the partial derivatives:

"N_F = (\frac{∂F}{∂x},\frac{∂F}{∂y},\frac{∂F}{∂z}) =(2,-3,6);"  

"N_G = (\frac{∂G}{∂x},\frac{∂G}{∂y},\frac{∂G}{∂z}) =(8x,-18y,72z)."


3. If the plain touches the conicoid, then their normal vectors at the point of contact must be collinear:

"N_G = k * N_F;"

"(8x,-18y,72z)=k*(2,-3,6);"

"x=\frac{k}{4}; y=\frac{k}{6}; z=\frac{k}{12}."


4. Put the values of variables to the equation of the plain and find the coordinates of the contact point:

"2*\frac{k}{4}-3*\frac{k}{6}+6*\frac{k}{12}-6=0;"

"\frac{k}{2}-\frac{k}{2}+\frac{k}{2}-6=0;"

"k=12;x=3;y=2;z=1;" M0(3,2,1).

This point lies also in the conicoid:

"4x\u00b2-9y\u00b2+36z\u00b2-36=4*3\u00b2-9*2\u00b2+36*1\u00b2-36=0."


5. Thus, M0 is a common point of surfaces, and normal vectors are collinear there, so the plain touches the conicoid at the point M0(3,2,1).